JEF meeting, April 18, 2025

From GlueXWiki
Jump to: navigation, search

Meeting time and location:

  • ECAL maintenance and performance during the run
    • ECAL scalers before the accelerator maintenance day [1]
    • Checking ECAL modules [2]
    • Issue with the FCAL HV module [3]
    • Flash ADC scalers during first beam [4]
    • Two modules fixed during the accelerator maintenance day [5]
      • ECAL scalers after maintenance day (production at 100 nA on 4.5x10^-4 Al radiator) [6]
    • Commissioning of the ECAL trigger [7] [8]
    • Issues with the readout crate during two first two production nights [9]
  • Calibrations
    • First look at Compton scattering events in the ECAL [10]
    • ECAL/FCAL energy and timing [11]
    • ECAL coverage with 2-arm Compton [12]
    • Pi0 calibration [13]
  • Monitoring
  • Restoring PS test setup [14]
    • Assembling PS prototype [15]

Minutes

Attendees: Sasha, Bill, Yuyao, Olga, Drew, Besha, Vlad, Payton, Axel, Jane, Igal, S.T.

  • ECAL HV has been updated using Compton events
    • Individual HV settings for 146 modules
    • Increased HV for most channels
    • Set energy scale for FADC: 2.8 MeV/ADC count
  • Maintenance day: used LMS to check detector
    • Found and fixed one hot channel and a couple other suspicious channels
  • Trigger timing between ECAL and FCAL has been adjusted
  • π0 calibrations under way
    • π0 peaks clearly seen in most channels, but not seen for ~10% of ECAL channels
    • Trying to understand missing peaks -- scalers look good, baselines are fine, not trigger problem...
    • New simplified peak fitting procedure
    • TOF not properly in time, needed for veto